[Hallc_running] [Revised Logentry] RC Update -- Saturday 08/09/25 [NO Meeting]

pdbforce at jlab.org pdbforce at jlab.org
Sat Aug 9 12:55:02 EDT 2025


Logentry Text:
--
<h2>Current Status</h2>
<ul>
  <li>Smooth production running on Dummy at 40 uA. This is the first Dummy cycle at z = 0.5 e-pi- setting (SHMS_theta = 11.07 deg). Next, we will move to the final LH2 cycle.</li>
</ul>

<h2>Key Updates</h2>
<ul>
  <li>Yesterday morning, MCC's attempt to swap the electronics card for the Hall C dump diffuser right IC did not solve the issue. They decided to mask it and run with only the Hall C dump diffuser left IC, with RadCon's authorization. Following that, they performed a quick BLA calibration and were able to give us 60 uA. However, at 60 uA the trip rate was a bit higher, so we requested 55 uA instead, which has been stable. We have been running at 55 uA max current since then.</li>
  <li>We finished the e-pi+ running last night around 11 PM. It then took about 1.5 hours to flip the SHMS magnet polarity for pi- running. A third of this time was spent troubleshooting the "cycle complete" indicator not turning green for SHMS Q2. Sam, the engineer on call, after consulting with Eleen, informed us that this is a known intermittent issue and safe to run in this condition. (I have requested Ed to include this information in the run plan to avoid possible time loss in the future.) <a href="https://logbooks.jlab.org/entry/4438177">4438177</a></li>
  <li>As soon as we were ready to take production data at z = 0.5 e-pi- setting (SHMS_theta = 11.07 deg) after the polarity flip, the Hall B solenoid tripped, taking down most of the HMS and SHMS magnets. It took about 3 hours to recover. Many thanks to Sam for helping us get back online as quickly as possible. <a href="https://logbooks.jlab.org/entry/4438193">4438193</a></li>
  <li>We resumed production on LH2 around 3:15 AM and finished the first cycle about an hour ago, around 11:45 AM. <a href="https://logbooks.jlab.org/entry/4438245">4438245</a></li>
  <li>Since the beginning of today's day shift, Hall B has tried to ramp up their solenoid twice, each time causing HMS Q1, Q3, and D interlocks. Each incident took about 30 minutes to recover from. Hall B RC, our engineers, and the Hall B team are aware of the issue and are trying to minimize such interruptions. <a href="https://logbooks.jlab.org/entry/4438393">4438393</a> <a href="https://logbooks.jlab.org/entry/4438486">4438486</a> </li>
</ul>

<h2>Links to Shift Summaries</h2>
<ul>
  <li>Day Shift (Friday) --- <a href="https://logbooks.jlab.org/entry/4437648">https://logbooks.jlab.org/entry/4437648</a></li>
  <li>Swing Shift (Friday) ---  <a href="https://logbooks.jlab.org/entry/4437937">https://logbooks.jlab.org/entry/4437937</a> </li>
  <li>Owl Shift (Saturday) --- <a href="https://logbooks.jlab.org/entry/4438171">https://logbooks.jlab.org/entry/4438171</a> </li>
</ul>

<h2>Run Plan for Today</h2>
<ul>
  <li>Continue taking production data @ z=0.5 e-pi- setting (SHMS_theta = 11.07 deg). </li>
</ul>

<h2>Summary of Statistics</h2>
<table border="1">
  <thead>
    <tr>
      <th>Target</th>
      <th>Goal</th>
      <th>Total cnts so far</th>
      <th>Cycles completed</th>
    </tr>
  </thead>
  <tbody>
    <tr>
      <td>LH2</td>
      <td>23k</td>
      <td>12973</td>
      <td>1/2</td>
    </tr>
    <tr>
      <td>Dummy</td>
      <td>4k</td>
      <td>0</td>
      <td>0/2</td>
    </tr>
  </tbody>
</table>
<a href="https://hallcweb.jlab.org/wiki/index.php?title=Kinematic_setting_summaries">Link to the kinematic setting summary table</a><br>

<br>
<b>We will not be meeting today. Please call the RC phone if you need to reach me.</b>


---

This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4438514
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20250809/22764030/attachment-0001.htm>


More information about the Hallc_running mailing list